OpenStudy (anonymous):

Part 1: Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided. Part 2: View and comment on the work of at least two other students. Try to choose students whose examples are different from the ones you created.

OpenStudy (anonymous):

first one Create a sample problem that has a binomial which IS a factor of the polynomial being divided, start with your polynomial in factored for, make up anything where one is a "binomial" say (x2+2x+1)(x−4) multiply out, and then you will be sure that x−4 is a factor of your result for the second one, you can take your answer for the first one and add some number to it. then you can be sure that x−4 is not a factor

OpenStudy (anonymous):

so what would that look like "?

OpenStudy (anonymous):

i have not much idea about binomials and polynomials i ask this from a friend so i dont have much idea how to actually make them

OpenStudy (anonymous):

http://www.myalgebra.com/algebra_solver.aspx try this it will help

OpenStudy (anonymous):

i m really really sorry sis that i cant help much

OpenStudy (accessdenied):

Do you know how to tell if a number like 13 is a factor of a number? Basically ,if you have a number with a factor of 13 (like 26, 2*13, 13 is a factor), then we should be able to divide that number by 13 and get no remainder. 26 / 13 = 2, 13 is a factor. 28 / 13 = 2 R: 2, with good reason since 28 = 2*2*2*3. no 13's there.

OpenStudy (anonymous):

what?

OpenStudy (accessdenied):

It's the same concept with polynomials: if we divide off a binomial from a polynomial and get no remainder, then the binomial would be a factor of that polynomial.

OpenStudy (anonymous):

so how would i answer this to make since to my teacher?

OpenStudy (accessdenied):

I could try my best to explain whatever you're unsure of. I'm just wondering what aspect of polynomials is confusing to you? Sorry, I don't want to just give you an answer, it doesn't really help you understand it. :(

OpenStudy (anonymous):

no like how am i suppose to answer this like am i suppose to give an example or explane? and if an example what should i write?